OpenStudy (anonymous):

How to simplify the expression: sin(4x)sin(3x)-cos(4x)cos(3x)

OpenStudy (goformit100):

"A warm Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (goformit100):

use compound angle formula for solving this particular answer.

zepdrix (zepdrix):

Here is the Cosine Angle Sum Identity: \[\Large \cos \alpha \cos \beta-\sin \alpha \sin \beta\quad=\quad \cos(\alpha + \beta)\] Hmm could we maybe alter our expression a little bit to fit this form?

OpenStudy (anonymous):

@zepdrix : oh...thnks, simplify is cos7x, right ?

OpenStudy (kainui):

Not quite, you need to look closer. @evahenny

OpenStudy (anonymous):

@Kainui : and then, what the answer for my problem ?

OpenStudy (kainui):

You tell me! Just write the problem you have and the identity right next to each other and compare them closely. You might notice something is slightly different.

OpenStudy (anonymous):

@Kainui : oh..thnks

OpenStudy (kainui):

=D

OpenStudy (anonymous):

@Kainui you are very smart :)